# A net theoretic approach to homotopy theory

### v1

@@ 3. Homotopy and fundamental groupoids in limit spaces
The previous example yields a non-topological space with trivial fundamental group [^11] For another easy example, recall that for a Tychonoff space $X$, $\operatorname{C}\left(X,\mathbb{R}\right)$ has a topology inducing the continuous convergence if and only if $X$ is locally compact [11]. So, $\left\langle\operatorname{C}\left(X,\mathbb{R}\right),\to_{c}\right\rangle$ is a non-topological limit vector space whenever $X$ is a Tychonoff and non-locally compact space, hence it is null-homotopic.. To get a non-topological space with non-trivial fundamental group, we adapt Example 1.3.2 in [1]. Let $\nu$ be the usual topology of $\mathbb{R}^{2}$ and $X\subseteq\mathbb{R}^{2}$ be a “lollipop”, i.e., the union of a line $L$ and a circle $C$ of the plane, such that $L\cap C=\{p\}$, as in Figure 1 below.
@@
The definition of $\lambda$ ensures $\left\langle X,\lambda\right\rangle\in\textsc{Lim}$. To see that $\left\langle X,\lambda\right\rangle\not\in\textsc{Top}$, notice that $\operatorname{adh}_{\lambda}\left(A\setminus C\right)=A\setminus\{p\}$ and $\operatorname{adh}_{\lambda}\left(A\setminus\{p\}\right)=A$, thus showing that the adherence operator is not idempotent. It is not hard to see that $\pi_{1}(X,p)=\pi_{1}(C,p)$: since the topological modification of $\left\langle X,L\right\rangle$ is the $\nu$-subspace topology inherited from $\mathbb{R}^{2}$, it follows that a continuous path $\gamma\colon I\to X$ such that $\gamma(0)=p$ cannot intercept $S$, i.e., there is no $t\in I$ such that $\gamma(t)\in S$, otherwise we would be able to obtain a sequence $\left\langle x_{n}\right\rangle_{n}$ in $I$ such that $x_{n}\to 0$ but $\gamma(x_{n})\in A\setminus C$ for every $n$.
@@ 4. The Seifert-Van Kampen Theorem for fundamental groupoids of limit spaces
The purpose of this section is to replace the word “topological space” with “limit space” in a reasonable way. To do that, all we have to do is to replace “open covers” with “convergence systems”, as defined in Section 2, and carry over the usual proof with the obvious adaptations.
@@ Acknowledgments
We would like to thank Germán Ferrer, Peter Wong and Weslem Silva for their insightful and helpful discussions on the subject.

### v2

@@ 3. Homotopy and fundamental groupoids in limit spaces
The previous example yields a non-topological space with trivial fundamental group [^11] For another easy example, recall that for a Tychonoff space $X$, $\operatorname{C}\left(X,\mathbb{R}\right)$ has a topology inducing the continuous convergence if and only if $X$ is locally compact [11]. So, $\left\langle\operatorname{C}\left(X,\mathbb{R}\right),\to_{c}\right\rangle$ is a non-topological limit vector space whenever $X$ is a Tychonoff and non-locally compact space (e.g., $X=\mathbb{Q}$), hence it is null-homotopic.. To get a non-topological space with non-trivial fundamental group, we adapt Example 1.3.2 in [1]. Let $\nu$ be the usual topology of $\mathbb{R}^{2}$ and $X\subseteq\mathbb{R}^{2}$ be a “lollipop”, i.e., the union of a line $L$ and a circle $C$ of the plane, such that $L\cap C=\{p\}$, as in Figure 1 below.
@@
The definition of $\lambda$ ensures $\left\langle X,\lambda\right\rangle\in\textsc{Lim}$. To see that $\left\langle X,\lambda\right\rangle\not\in\textsc{Top}$, notice that $\operatorname{adh}_{\lambda}\left(S\setminus D\right)=S\setminus\{p\}$ and $\operatorname{adh}_{\lambda}\left(S\setminus\{p\}\right)=L$, thus showing that the adherence operator is not idempotent. It is not hard to see that $\pi_{1}(X,p)=\pi_{1}(C,p)$: since the topological modification of $\left\langle X,L\right\rangle$ is the $\nu$-subspace topology inherited from $\mathbb{R}^{2}$, it follows that a continuous path $\gamma\colon I\to X$ such that $\gamma(0)=p$ cannot intercept $S$, i.e., there is no $t\in I$ such that $\gamma(t)\in S$, otherwise we would be able to obtain a sequence $\left\langle x_{n}\right\rangle_{n}$ in $I$ such that $x_{n}\to 0$ but $\gamma(x_{n})\in S\setminus D$ for every $n$, preventing $\gamma(x_{n})\to_{\lambda}\gamma(0)=p$ to happen.
Remark 3.7.
In a private communication, Antonio Rieser kindly pointed out that the existence of fundamental group(oid) for limit spaces can also be derived from the cofibration structure he presented in [34]. We emphasize that our main purpose here is to illustrate how nets can simplify the treatment of convergence spaces.
@@ 4. The Seifert-Van Kampen Theorem for fundamental groupoids of limit spaces
The purpose of this section is to replace the word “topological space” with “limit space” in a reasonable way. All we have to do is to replace “open covers” with “convergence systems”, as defined in Section 2, and carry over the usual proof with the obvious adaptations.
@@ 5. Further remarks and comments
In this regard, we should also mention that the category Lim may be overly broad for further inquiries. On one hand, compactness is central to any considerations in Topology, and for this, the category of pseudotopological spaces appears as a natural setting, being the most suitable environment for handling compactness (cf. [10, 11]). On the other hand, as shown by Rieser in [34], pseudotopological suspensions coincide with their topological counterparts when applied to spheres, ensuring that many homotopy-theoretic results from topological spaces carry over to pseudotopological spaces. Thus, our choice of Lim in this work was motivated by its minimal hypotheses needed for our current purposes.
Remark 5.1 (On pseudotopologies and nets).
Recall that a convergence space $\left\langle X,L\right\rangle$ is pseudotopological if
$L(\mathcal{F})=\bigcap_{\mathfrak{u}\in\beta(\mathcal{F})}L(\mathfrak{u})$
holds for every proper filter $\mathcal{F}$ on $X$, where $\beta(\mathcal{F})$ stands for the set of ultrafilters on $X$ containing $\mathcal{F}$ (cf. [1]). In this sense, ultrafilters can be easily replaced by ultranets, which are precisely those nets whose induced filters are ultrafilters [^13] Also called universal nets. Equivalently, $\varphi\in\textsc{NETS}\left(X\right)$ is an ultranet if, and only if, $\varphi$ is a subnet of every $\psi\in\textsc{NETS}\left(X\right)$ which is a subnet of $\varphi$. (cf. [35]). Alternatively, it is not hard to see that $\left\langle X,L\right\rangle$ is pseudotopological if and only for every net $\varphi\in\textsc{NETS}\left(X\right)$ and every point $x\in X$, $\varphi\to_{L}x$ whenever every subnet $\psi$ of $\varphi$ has a further subnet $\rho$ such that $\rho\to_{L}x$, which is the exact translation of the topological fact motivating the definition of pseudotopologies to begin with (cf. [25, Section 5.1.4]).
@@ Acknowledgments
We would like to thank Germán Ferrer, Peter Wong, and Weslem Silva for their insightful and helpful discussions on the subject. We also extend our gratitude to Antonio Rieser for his thoughtful observations.